# GENPACK: KPI-Guided Multi-Objective Genetic Algorithm for Industrial 3D Bin Packing

### v1

@@ 1. Introduction
Constructive heuristics such as Bottom-Left (chazelle1983bottomleft, ), Extreme Point (crainic2008extreme, ), or MaxRects (jylanki2010maxrects, ) remain widely used for their efficiency, particularly on large instances (bischoff1990three, ). However, they rely heavily on simplified placement rules that often yield arrangements failing to satisfy industrial safety and stability constraints. Metaheuristics, and in particular genetic algorithms (GAs), provide greater flexibility by exploring large solution spaces and accommodating multiple objectives (jakobs1996genetic, ; wu2010three, ). However, pure GA approaches also face limitations: common encodings are not well suited to the packing structure, genetic operators often fail to preserve feasibility, and fitness functions often fail to incorporate practical indicators of industrial quality (bortfeldt2001hybrid, ; egeblad2009placement, ; bortfeldt2013constraints, ). Moreover, scalability to real-world order sizes at industrial pallet dimensions, such as Euro-pallets and roll containers, remains a critical challenge where many state-of-the-art methods still struggle (wu2010three, ; egeblad2009placement, ).
Recent work has introduced hybrid methods that combine heuristics with metaheuristic refinement. For example, Ananno and Ribeiro (ananno2024multiheuristic, ) proposed a two-stage framework in which constructive heuristics generate layers and blocks, and a GA arranges the remaining residual items under stability and support constraints. While such approaches demonstrate feasibility, they primarily enforce industrial requirements as hard constraints and focus evaluation on compactness and utilization (bortfeldt2001hybrid, ; kang2012hybrid, ; egeblad2009placement, ). They do not systematically optimize across multiple industrial key performance indicators (KPIs), and runtime scalability remains problematic for highly heterogeneous orders (egeblad2009placement, ; wu2010three, ).
@@ 2.2. Metaheuristics
The first application of genetic algorithms (GAs) to packing was introduced by Jakobs (jakobs1996genetic, ), with subsequent work focusing on hybridizing GAs with constructive methods (wu2010three, ; kang2012hybrid, ). Later studies showed that problem-specific encodings and tailored operators perform better than generic representations (karabulut2004three, ), and that combining GAs with local search often yields state-of-the-art results (ceschia2013metaheuristic, ). Despite these advances, many GA implementations still rely on simplified fitness functions that fail to capture industrial constraints and continue to face runtime scalability issues, particularly for heterogeneous orders.
@@ 4.2. Multi-KPI Fitness Function
  • RelDen: relative density below the upper envelope, capturing void reduction and local compactness. While compaction is a common heuristic principle in 3D-BPP algorithms (crainic2008extreme, ; egeblad2009placement, ), this envelope-based voxel KPI is specific to this work. Using a voxel grid of spatial resolution $r$ discretizing the pallet base $(L,W)$ and height $H^{\ast}$, let $\mathcal{G}[x,y,z]\in\{0,1\}$ denote occupancy (1 if voxel is filled). Define the upper envelope
@@ 4.5. Parameters
We use a population size of 100, up to 50 generations, an elite rate of 10%, and adaptive $P_{c}$ and $P_{m}$ for crossover and mutation, respectively, as described above. Fitness values are clipped to a small positive lower bound for numerical stability. All parameter values fall within ranges commonly adopted in GA-based three-dimensional bin packing and container loading literature (jakobs1996genetic, ; karabulut2004three, ; wu2010three, ; kang2012hybrid, ; bortfeldt2013constraints, ; ananno2024multiheuristic, ). Sensitivity experiments within typical ranges ($P_{c}\in[0.4,0.6]$, $P_{m}\in[0.2,0.4]$, population =50=50200200) confirmed similar convergence behavior. The adaptive scheduling of $P_{c}$ and $P_{m}$ gradually increases exploitation and reduces randomness toward later generations, a standard mechanism for stabilizing GA search (eiben1999parameter, ).

### v2

@@ 1. Introduction
Constructive heuristics such as Bottom-Left (chazelle1983bottomleft, ), Extreme Point (crainic2008extreme, ), or MaxRects (jylanki2010maxrects, ) remain widely used for their efficiency, particularly on large instances (bischoff1990three, ). However, they rely heavily on simplified placement rules that often yield arrangements failing to satisfy industrial safety and stability constraints. Metaheuristics, and in particular genetic algorithms (GAs), provide greater flexibility by exploring large solution spaces and accommodating multiple objectives (jakobs1996genetic, ; wu2010three, ). However, pure GA approaches also face limitations: common encodings are not well suited to the packing structure, genetic operators often fail to preserve feasibility, and fitness functions often fail to incorporate practical indicators of industrial quality (bortfeldt2001hybrid, ; egeblad2009heuristic, ; bortfeldt2013constraints, ). Moreover, scalability to real-world order sizes at industrial pallet dimensions, such as Euro-pallets and roll containers, remains a critical challenge where many state-of-the-art methods still struggle (wu2010three, ; egeblad2009heuristic, ).
Recent work has introduced hybrid methods that combine heuristics with metaheuristic refinement. For example, Ananno and Ribeiro (ananno2024multiheuristic, ) proposed a two-stage framework in which constructive heuristics generate layers and blocks, and a GA arranges the remaining residual items under stability and support constraints. While such approaches demonstrate feasibility, they primarily enforce industrial requirements as hard constraints and focus evaluation on compactness and utilization (bortfeldt2001hybrid, ; kang2012hybrid, ; egeblad2009heuristic, ). They do not systematically optimize across multiple industrial key performance indicators (KPIs), and runtime scalability remains problematic for highly heterogeneous orders (egeblad2009heuristic, ; wu2010three, ).
@@ 2.2. Metaheuristics
The first application of GAs to packing was introduced by Jakobs (jakobs1996genetic, ), with subsequent work focusing on hybridizing GAs with constructive methods (wu2010three, ; kang2012hybrid, ). Later studies demonstrated that problem-specific encodings and tailored operators consistently outperform generic representations (gehring1997genetic, ; bortfeldt2001hybrid, ), and that combining GAs with local search often yields state-of-the-art results (ceschia2013metaheuristic, ). Despite these advances, many GA implementations still rely on simplified fitness functions that fail to capture industrial constraints and continue to face runtime scalability issues, particularly for heterogeneous orders.
@@ 4.2. Multi-KPI Fitness Function
  • RelDen: relative density below the upper envelope, capturing void reduction and local compactness. While compaction is a common heuristic principle in 3D-BPP algorithms (crainic2008extreme, ; egeblad2009heuristic, ), this envelope-based voxel KPI is specific to this work. Using a voxel grid of spatial resolution $r$ discretizing the pallet base $(L,W)$ and height $H^{\ast}$, let $\mathcal{G}[x,y,z]\in\{0,1\}$ denote occupancy (1 if voxel is filled). Define the upper envelope
@@ 4.5. Parameters
We use a population size of 100, up to 50 generations, an elite rate of 10%, and adaptive $P_{c}$ and $P_{m}$ for crossover and mutation, respectively, as described above. Fitness values are clipped to a small positive lower bound for numerical stability. All parameter values fall within ranges commonly adopted in GA-based three-dimensional bin packing and container loading literature (jakobs1996genetic, ; wu2010three, ; kang2012hybrid, ; bortfeldt2013constraints, ; ananno2024multiheuristic, ). Sensitivity experiments within typical ranges ($P_{c}\in[0.4,0.6]$, $P_{m}\in[0.2,0.4]$, population =50=50200200) confirmed similar convergence behavior. The adaptive scheduling of $P_{c}$ and $P_{m}$ gradually increases exploitation and reduces randomness toward later generations, a standard mechanism for stabilizing GA search (eiben1999parameter, ).